# A New Dataset and Performance Benchmark for Real-time Spacecraft Segmentation in Onboard Flight Computers

### v1

@@ Abstract
Spacecraft deployed in outer space are routinely subjected to various forms of damage due to exposure to hazardous environments. In addition, there are significant risks to the subsequent process of in-space repairs through human extravehicular activity or robotic manipulation, incurring substantial operational costs. Recent developments in image segmentation could enable the development of reliable and cost-effective autonomous inspection systems. While these models often require large amounts of training data to achieve satisfactory results, publicly available annotated spacecraft segmentation data are very scarce. Here, we present a new dataset of nearly 64k annotated spacecraft images that was created using real spacecraft models, superimposed on a mixture of real and synthetic backgrounds generated using NASA’s TTALOS pipeline. To mimic camera distortions and noise in real-world image acquisition, we also added different types of noise and distortion to the images. Finally, we finetuned YOLOv8 and YOLOv11 segmentation models to generate performance benchmarks for the dataset under well-defined hardware and inference time constraints to mimic real-world image segmentation challenges for real-time onboard applications in space on NASA’s inspector spacecraft. The resulting models, when tested under these constraints, achieved a Dice score of 0.92, Hausdorff distance of 0.69, and an inference time of about 0.5 second. The dataset and models for performance benchmark are available at https://github.com/RiceD2KLab/SWiM.
Index Terms:
spacecraft, segmentation, dataset, YOLO
@@ I Introduction
The computational constraints of orbital hardware necessitate architectures balancing inference time and precision. We adapt the You Only Look Once (YOLO) framework[3], renowned for its real-time inference capabilities, for spacecraft segmentation tasks. While other CNN-based object detection and segmentation models, such as Mask-RCNN [4] excel in similar tasks, their computational requirements and inference time often exceed the capabilities of resource-constrained flight-grade hardware. The unified detection-segmentation approach in recent variants of YOLO, such as YOLOv8 [5] and YOLOv11 [6], enables efficient processing on our target platform, a Universal Platform (UP) Board Intel Celeron N3350 CPU with 4GB RAM, which is used by NASA for on-board flight computers in inspector spacecraft.
@@
  • Our problem formulation codifies real-world deployment requirements through quantifiable constraints on hardware (inference on CPU with $<$4 GB RAM only) and inference latency ($<$0.95s).
@@ III-B1 Synthetic Image Generation Methods
Method 1 (Superimposing segmentations with backgrounds for Spacecrafts dataset)
@@
Method 2 (TTALOS Pipeline with Stable Diffusion for PoseBowl dataset)
@@
For generating foreground spacecraft images, we used NASA’s Toolset for Training and Labeling in an Optical Simulator (TTALOS) pipeline and Blender to render 2D projections of unique 3D spacecraft models [7]. In particular, we leveraged $15$ unique 3D spacecraft renderings in the training, $5$ in the validation, and $5$ in the test partitions of SWiM. Each rendering featured distinct designs and colors. The 3D spacecraft renderings were generated with randomized poses, lighting conditions, and scales, initially placed on transparent backgrounds. Each spacecraft image was then superimposed onto a unique background generated using Stable Diffusion, resulting in composite images with associated target segmentation masks.
@@
III-C SWiM Dataset Versions
@@ III-C1 Baseline
III-C2 Augmented
@@
Figure 4: Sample synthetic images in the augmented dataset.
@@ IV Models for Performance Benchmarks
While each YOLO model is available in multiple variants, depending on size and inference time, we chose the nano versions of the models due to the computational and operational constraints imposed by the target deployment environment: NASA’s Intel UP Board. This board, being a compact single-board computer, offers limited hardware capabilities, including the absence of a GPU, only 4 GB of RAM memory, and 4 CPU cores [20]. Consequently, the segmentation algorithm must not only achieve high accuracy but also perform inference in real-time, ideally under 0.95s per image, to support spacecraft inspection during proximity operations in space [1].
@@ V-A2 Hausdorff Distance :
where $X$ is the predicted set of pixels, $Y$ is the ground truth set of pixels, and $d(x,y)$ is the Euclidean distance between points x and y. We note here that Hausdorff Distance provides insight into the worst-case segmentation error and is particularly useful for evaluating the accuracy of object boundaries in segmentation tasks. It complements the Dice coefficient by focusing on the spatial distribution of segmentation errors rather than simple volumetric overlap.
@@ V-B Experiments
Further, to simulate the Universal Platform (UP) Board Intel Celeron N3350 CPU with 4GB RAM, a docker image and container were created, and limited to 3 GB RAM and 3 CPU cores, leaving aside the remaining hardware for contingency. This extreme buffer design choice ensured that the model was tested under tighter constraints for inference time.
@@ V-C1 Performance Benchmark for Segmentation
TABLE II: Performance Benchmark using YOLO v8n and YOLO v11n trained on baseline and augmented versions of SWiM.
Model SWiM Version Avg. Dice Score Avg. Hausdorff
YOLO v8n Baseline 0.92 0.69
YOLO v8n Augmented 0.92 0.88
YOLO v11n Baseline 0.92 1.07
YOLO v11n Augmented 0.92 0.74
Table II presents the test performance of the two versions of the YOLO segmentation model, YOLOv8 nano and YOLOv11 nano trained on the baseline and augmented training sets using Dice coefficient and Hausdorff Distance. We note that all the models achieved a high average Dice score of $0.92$ on the test dataset, indicating that the masks are generally very accurate in identifying the bulk area of the spacecraft. All 4 models have a very low average Hausdorff distance between $0.69$-$1.07$, indicating that the models are generally excellent at detecting the complex boundaries of various spacecraft models.
Overall, YOLO v8 and v11 both achieve excellent performance in detecting the bulk of the spacecraft, as well as the boundaries of most spacecraft. We note that using the augmented version of SWiM does not necessarily have a significant impact on the performance of YOLO. We note that the models fulfill the primary performance objective of accurately detecting spacecraft without excessive false positives, though they may miss a few pixels due to their conservative approach.
@@ V-C2 Inference Time and Model Size
TABLE III: Benchmark Inference Time for Models.
Model SWiM Version Inf. (ms) FPS Size (MB)
YOLO v8n Baseline 443.90 2.25 12.7
YOLO v8n Augmented 586.78 1.87 12.7
YOLO v11n Baseline 533.62 1.7 11.1
YOLO v11n Augmented 446.93 2.24 11.1
YOLOv8 nano was slightly larger and had a marginally higher average inference time of 515.34 ms across baseline and augmented data (compared to 490.275 ms for YOLOv11 nano). YOLOv8 nano showed a substantial increase in inference time from the baseline to the augmented dataset. However, the inference time of YOLOv11 nano improved for the augmented dataset. However, the differences are not high enough to be conclusive and may be attributed to resource allocation processes during the simulation. Overall, both models have an approximate inference time of 500 ms which is nearly half the time constraint of 0.95s.
@@ VI Discussion
In addition, we define the constraints for model development considering deployment on resource-constrained onboard hardware in space flight and inference time limits for real-time operation (based on the current operating specifications of NASA). Using these constraints, we obtain and report benchmark performance results using YOLOv8 nano and YOLOv11 nano models. We note that YOLOv11 nano achieves comparable inference speed to YOLOv8 nano while being more storage-efficient, aligning with the project’s objectives for efficient, edge-compatible deployment. Both models successfully met all technical constraints of on-board deployment by NASA, operating efficiently within a constraint-specific Docker container ($<$4GB RAM with 4 cores and CPU-only inference) and achieving an inference time under 0.95s per image, as required for real-time performance in autonomous spacecraft inspection.

### v2

# A New Dataset and Performance Benchmark for Real-time Spacecraft Segmentation in Onboard Computers

@@ Abstract
Spacecraft deployed in outer space are routinely subjected to various forms of damage due to exposure to hazardous environments. In addition, there are significant risks to the subsequent process of in-space repairs through human extravehicular activity or robotic manipulation, incurring substantial operational costs. Recent developments in image segmentation could enable the development of reliable and cost-effective autonomous inspection systems. While these models often require large amounts of training data to achieve satisfactory results, publicly available annotated spacecraft segmentation data are very scarce. Here, we present a new dataset of nearly 64k annotated spacecraft images that was created using real spacecraft models, superimposed on a mixture of real and synthetic backgrounds generated using NASA’s TTALOS pipeline. To mimic camera distortions and noise in real-world image acquisition, we also added different types of noise and distortion to the images. Our dataset includes images with several real-world challenges, including noise, camera distortions, glare, varying lighting conditions, varying field of view, partial spacecraft visibility, brightly-lit city backgrounds, densely patterned and confounding backgrounds, aurora borealis, and a wide variety of spacecraft geometries. Finally, we finetuned YOLOv8 and YOLOv11 models for spacecraft segmentation to generate performance benchmarks for the dataset under well-defined hardware and inference time constraints to mimic real-world image segmentation challenges for real-time onboard applications in space on NASA’s inspector spacecraft. The resulting models, when tested under these constraints, achieved a Dice score of 0.92, Hausdorff distance of 0.69, and an inference time of about 0.5 second. The dataset and models for performance benchmark are available at https://github.com/RiceD2KLab/SWiM.
@@ I Introduction
While large datasets, such as SPEED+ [16], are available for pose estimation of spacecraft, these datasets do not include the pixel-level segmentation masks of spacecraft for training models. Furthermore, these datasets have limited diversity of spacecraft images. For example, SPEED+ consists of images of only a single spacecraft (Tango) generated using an OpenGL-based emulator (SLAB), whereas our dataset includes multiple spacecraft generated via the TTALOS pipeline.
The computational constraints of orbital hardware necessitate architectures balancing inference time and precision. We adapt the You Only Look Once (YOLO) framework[18], renowned for its real-time inference capabilities, for spacecraft segmentation tasks. While other CNN-based object detection and segmentation models, such as Mask-RCNN [9] excel in similar tasks, their computational requirements and inference time often exceed the capabilities of resource-constrained flight-grade hardware. The unified detection-segmentation approach in recent variants of YOLO, such as YOLOv8 [10] and YOLOv11 [11], enables efficient processing on our target onboard flight computer with 4GB RAM, which is used by NASA for on-board flight computers in inspector spacecraft.
@@
  • Our problem formulation codifies real-world deployment requirements through quantifiable constraints on hardware (inference on CPU with <<4 GB RAM only) and inference latency (<<0.95s).
@@ III-A1 PoseBowl
PoseBowl was created primarily using NASA’s Toolset for Training and Labeling in an Optical Simulator (TTALOS). The main workflow of TTALOS leverages high-fidelity 3D spacecraft models to produce annotated, photorealistic training data for computer vision applications. These 3D spacecraft models originate primarily from NASA, with a significant subset contributed by European Space Agency (ESA) for broader coverage of space missions, including spacecraft models such as Rosetta, Aqua, Aura, Mars Odyssey, Cassini, and Voyager.
@@ III-B1 Synthetic Image Generation Methods
Our synthetic dataset generation included three primary aspects:
  • Scene Construction and Backgrounds: The pipeline uses real satellite-captured images as backgrounds for scene rendering, creating high realism in composite images. However, synthetically generated backgrounds are also employed to introduce noise and randomness, further enhancing the diversity and robustness of the dataset. This enables superimposition of realistic 3D spacecraft over both authentic and artificial 2D backgrounds, creating variance desirable for training robust machine learning models.
  • Spacecraft Positioning, Orientation, and Lighting: The rendering script dynamically positions and rotates the 3D spacecraft models within each scene. This is achieved through random transformations, ensuring coverage of a wide range of visual perspectives. The lighting system is a key feature that uses two modes: in one mode, lighting is synchronized with parameters extracted from the real background image metadata (e.g., sun position and intensity); in the other mode, lighting energy and position are randomized, simulating diverse space illumination conditions. This results in images with spacecraft illuminated and shadowed in a number of realistic combinations.
  • Augmentation and Compositing Effects: The pipeline supports several image augmentations for added realism and domain generalization:
    1. 1.
      Glare: Different types such as fog glow, streaks, simple star, and ghost glare simulate lens artifacts and environmental glare.
    2. 2.
      Blur: Adjustable via blending nodes, mimicking motion blur or camera focus artifacts.
    3. 3.
      Exposure: Randomized exposure settings create effects ranging from underexposure to bright highlight clipping.
    These compositing augmentations are controlled by a function which randomizes parameters for each rendered image. Camera and environmental effects can be toggled and tuned according to the desired variance in the generated dataset.
  • Annotation and Data Export: For every rendered image, the pipeline saves both the image file and corresponding bounding box coordinates that locate the spacecraft within the image.
@@
III-B2 Method 1 (Superimposing segmentations with backgrounds for Spacecrafts dataset)
@@
III-B3 Method 2 (TTALOS Pipeline with Stable Diffusion for PoseBowl dataset)
@@
For generating foreground spacecraft images, we used NASA’s Toolset for Training and Labeling in an Optical Simulator (TTALOS) pipeline and Blender to render 2D projections of unique 3D spacecraft models [2]. . In addition to NASA spacecraft models, our dataset incorporated a wider selection of spacecraft from multiple agencies and organizations, including the European Space Agency (ESA), Satellite Communication Service (SES) from Luxembourg, and Johns Hopkins University Applied Physics Laboratory (APL). In particular, we leveraged $15$ unique 3D spacecraft renderings in the training, $5$ in the validation, and $5$ in the test partitions of SWiM. Each rendering featured distinct designs and colors. The 3D spacecraft renderings were generated with randomized poses, lighting conditions, and scales, initially placed on transparent backgrounds. Each spacecraft image was then superimposed onto a unique background generated using Stable Diffusion, resulting in composite images with associated target segmentation masks.
To ensure unbiased and reliable evaluation of model generalization, the spacecraft models included in the training, validation, and test partitions of the SWiM dataset are mostly non-overlapping. This design largely prevents the models from overfitting by memorizing specific spacecraft features seen during training, instead encouraging generalization to novel spacecraft designs, colors, and structures. However, a few spacecraft models are deliberately included in both training and validation/test sets as controlled inclusions, which allow for domain adaptation analysis or robustness checks across different lighting and background conditions.
The spacecraft models for each partition are as follows:
  1. 1.
    Training Set (15 models): Rosetta, Aqua, Astra, Aura, Deep Space 1, Mars Odyssey, Messenger, TDRS, Ulysses, Pioneer, Juno, MGS, Galileo, Near, and Ibex.
  2. 2.
    Validation (5 models): Grace, NPP, EPOXI, Dawn, Acrimsat
  3. 3.
    Test (5 models): Voyager, Themis, Cassini 66, Stardust, MRO
This structured partitioning strategy enables a robust training process while facilitating realistic evaluation scenarios that test the model’s ability to perform well on unseen spacecraft configurations.
@@
Beyond reusing satellite images, we greatly increase scene complexity by generating custom backgrounds using the diffusion model. This approach inserts rare, unexpected, or extreme conditions that may be underrepresented or absent in satellite-captured images:
  • Carefully crafted text prompts are used to instruct the diffusion model to create highly varied, high-resolution space backgrounds. These range from faint starlit voids, dense star fields, blurred planetary fly-by, dramatic sun glares, atmospheric auroras, to night-time city lights.
  • Sensor and atmospheric artifacts, such as chromatic aberration, lens flare, pixelation, and simulated sensor noise, are described in the prompt and infused into the output, resulting in backgrounds that challenge and generalize the training data far beyond ordinary real satellite imagery.
@@
III-C Capturing Real Imaging Complexity
Overall, our pipeline is specifically designed to fully capture the unpredictability and complexity of real orbital imaging conditions by integrating both natural and artificially generated variations at every stage of data creation. We capture real imaging complexity through the following means:
  1. 1.
    Lighting Variability: We randomize both the intensity and spatial position of the primary light source, mimicking rapid changes in solar illumination and shadowing experienced in orbit.
  2. 2.
    Camera Effects: Augmentations such as random blur (to simulate motion or out-of-focus artifacts), exposure variation (for under/overexposed scenes), and glare (from sun angles, sensor mechanics, or rapid light transitions) are systematically applied. These imitate the unpredictable optical effects captured by real orbital cameras. Diverse Backgrounds: By default, scenes use real satellite images; we further add complexity to this by adding synthetic backgrounds.
III-D Diversity of Field of View and Occlusion
PoseBowl features images in which spacecraft typically occupy a small portion of the overall image area; over 85% of images contain spacecraft that cover less than 10% of the total image area. In comparison, the Spacecraft dataset is characterized by a more uniform distribution of spacecraft size within images; spacecraft spanning a broader range, from 10% up to 80% of the image area.
Importantly, the baseline Posebowl dataset is approximately eight times larger than the Spacecraft dataset, ensuring that, even after merging, the combined dataset will still contain a substantial number of images with small field-of-view (FoV) spacecraft. This large sample size closely mirrors real-world orbital imaging, where spacecraft are often distant or only partially captured within each frame, preserving a crucial aspect of operational realism. Nonetheless, the inclusion of a broader range of spacecraft sizes from the Spacecraft dataset is highly beneficial – it introduces valuable variability that supports robust model generalization and improves segmentation performance across both small and large FoVs, ultimately aligning the data distribution more closely with the full spectrum of scenarios encountered in practice.
Moreover, the presence of occlusion in our dataset allows segmentation models to learn how to identify and delineate spacecraft even when they are partially blocked by other objects or image artifacts, closely reflecting real-world orbital imaging scenarios. This strengthens the model’s ability to recognize the full extent of objects within cluttered or challenging visual environments,
III-E SWiM Dataset Versions
@@ III-E1 Baseline
III-E2 Augmented
Figure 4: Sample synthetic images in the augmented dataset. Note the diversity in background and the real-world challenges, such as glare and confounding background.
@@ III-E2 Augmented
In this work, we define “synthetically generated images” as composite samples created by generating virtual backgrounds via Stable Diffusion and integrating them into the TTALOS pipeline to superimpose 3D spacecraft models. These synthetic composites are exclusively allocated to the training partition to enhance environmental diversity. Conversely, the validation and test partitions consist entirely of images featuring real orbital backgrounds derived from the original PoseBowl dataset. This stratification ensures that while the model is exposed to high-variance synthetic domains during training, it is benchmarked strictly against photorealistic imagery representative of actual operational conditions.
@@ IV Models for Performance Benchmarks
While each YOLO model is available in multiple variants, depending on size and inference time, we chose the nano versions of the models due to the computational and operational constraints imposed by a resource-constrained target deployment environment. The hardware, being a compact single-board computer, offers limited hardware capabilities, including the absence of a GPU, only 4 GB of RAM memory, and 4 CPU cores [20]. Consequently, the segmentation algorithm must not only achieve high accuracy but also perform inference in real-time, ideally under 0.95s per image, to support spacecraft inspection during proximity operations in space [5].
@@ V-A2 Hausdorff Distance :
where $X$ is the predicted set of pixels, $Y$ is the ground truth set of pixels, and $d(x,y)$ is the Euclidean distance between points $x$ and $y$. We note here that Hausdorff Distance provides insight into the worst-case segmentation error and is particularly useful for evaluating the accuracy of object boundaries in segmentation tasks. In particular, Hausdorff Distance evaluates the boundaries of the segmentation masks and is not a simple region-based calculation that is agnostic to the boundaries of masks, such as Dice Coefficient. Thus, it complements the Dice coefficient by focusing on the spatial boundaries of segmentation masks rather than a simple volumetric overlap.
@@ V-B Experiments
Further, to simulate a typical resource-constrained onboard flight computer with 4GB RAM, a docker image and container were created, and limited to 3 GB RAM and 3 CPU cores, leaving aside the remaining hardware for contingency. This extreme buffer design choice ensured that the model was tested under tighter constraints for inference time.
@@ V-C1 Performance Benchmark for Segmentation
TABLE II: Performance Benchmark using YOLO v8n and YOLO v11n trained on baseline and augmented versions of SWiM. (Base stands for Baseline; Aug stands for Augmented)
Model
SWiM
Version
Average
Dice Score
Average
Hausdorff
YOLO v8n Base 0.9292 0.6895
YOLO v8n Aug 0.9269 0.8857
YOLO v11n Base 0.9287 1.0781
YOLO v11n Aug 0.9274 0.7476
Table II presents the test performance of the two versions of the YOLO segmentation model, YOLOv8 nano and YOLOv11 nano trained on the baseline and augmented training sets using Dice coefficient and Hausdorff Distance. We note that all the models achieved a high average Dice score of $0.92$ on the test dataset, indicating that the masks are generally very accurate in identifying the bulk area of the spacecraft. All 4 models have a very low average Hausdorff distance between 0.690.69-$1.07$, indicating that the models are generally excellent at detecting the complex boundaries of various spacecraft models.
Overall, YOLO v8 and v11 both achieve excellent performance in detecting the bulk of the spacecraft, as well as the boundaries of most spacecraft. We note that using the augmented version of SWiM does not necessarily have a significant impact on the performance of YOLO.
The consistent accuracy for both YOLO v8 and v11 verifies that both the datasets result in highly accurate models across different YOLO architectures. The difference in Dice coefficient between baseline and augmented datasets is of the order of $10^{-4}$, which indicates that the YOLO models are robust to the additional noise added in the augmented dataset. As this noise mimics several real-world scenarios such as camera errors and lighting issues, this robustness is very promising.
Based on the success of YOLO architectures on both datasets, we propose that the baseline and augmented datasets provide two options in terms of size for future work using other model architectures.
@@ V-C2 Inference Time and Model Size
TABLE III: Benchmark Inference Time for Models. (Base stands for Baseline; Aug stands for Augmented)
Model
SWiM
Version
Inf. (ms) FPS Size (MB)
YOLO v8n Base 443.90 2.25 12.7
YOLO v8n Aug 586.78 1.87 12.7
YOLO v11n Base 533.62 1.7 11.1
YOLO v11n Aug 446.93 2.24 11.1
The models successfully achieved near real-time inference speeds, with an average inference time of 500 ms, which is nearly half the goal time constraint of 0.95s. Though the YOLO v11 claims a more optimized model architecture, we observed only minor differences in the performances of v8 vs v11. These differences can be attributed to resource allocation processes during the simulation. Thus, both the v8 and v11 prove to be excellent choices based on performance as well as accuracy.
@@ VI Discussion
In addition, we define the constraints for model development considering deployment on resource-constrained onboard hardware in space flight and inference time limits for real-time operation (based on the current operating specifications of NASA). Using these constraints, we obtain and report benchmark performance results using YOLOv8 nano and YOLOv11 nano models. We note that YOLOv11 nano achieves comparable inference speed to YOLOv8 nano while being more storage-efficient, aligning with the project’s objectives for efficient, edge-compatible deployment. Both models successfully met all technical constraints of on-board deployment by NASA, operating efficiently within a constraint-specific Docker container (<<4GB RAM with 4 cores and CPU-only inference) and achieving an inference time under 0.95s per image, as required for real-time performance in autonomous spacecraft inspection.
@@
We also note certain limitations of our dataset. For segmentation mask generation, SAM2 was fine-tuned on a small sample of PoseBowl, in which the majority of the images had a field of view (FOV) less than 0.1 of the area of the image. Consequently, the SAM2 masks are likely to be the most accurate for similar images. For the 509 images with an FOV of over 0.5, we found that 98 presented inaccuracies that we manually corrected. The inaccurate masks produced by SAM2 primarily exhibited two common themes: misclassification of background pixels and edge noise. In some cases, the mask includes pixels from the background, treating them as part of the spacecraft structure. In others, the edges of the spacecraft are poorly defined, leading to imprecise segmentation boundaries.